a computer has two processors. the probability that processor 1 works is ¾ and the probability that processor 2 works is 7/8. the computer will function if processor 1 works or 2 works or both works. find probability thta processor 1 works given thta computeris functional

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Student
 
It can be solved in the following way>> 

Let denote processor 1 by "I" and processor2 by "II" , Computer by "C" 
P(I)=3/4    => P(I')=1/4 
P(II)=7/8   => P(II')=1/8 

NOW, P(C)=processor one works or processor two works or both processors work. 
  
=>P(C)=3/4* 1/8 + 7/8 * 1/4 + 3/4 * 7/8 
        
          =31/32 
Also, P(computer is funtional given processor1 is working)= 3/4* 1/8 + 3/4*7/8 
                                                                             =24/32 

Therefore, 

P(processor1 is working given computer is functional)= P(computer is funtional given processor1 is working) 
                                                                        --------------------------------------------------- 
                                                                                     P(C) 

                                                          = 24/32 
                                                           ------- 
                                                             31/32 

                                                           = 24/31
 
Regards
Arun (askIITians forum expert)